# Modeling, Inference, and Prediction in Mobility-Based Compartmental Models for Epidemiology

### v1

@@ Abstract
Classical compartmental models in epidemiology often struggle to accurately capture real-world dynamics due to their inability to address the inherent heterogeneity of populations. In this paper, we introduce a novel approach that incorporates heterogeneity through a mobility variable, transforming the traditional ODE system into a system of integro-differential equations that describe the dynamics of population densities across different compartments. Our results show that, for the same basic reproduction number, our mobility-based model predicts a smaller final pandemic size compared to classic compartmental models, whose population densities are represented as Dirac delta functions in our density-based framework. This addresses the overestimation issue common in many classical models. Additionally, we demonstrate that the time series of the infected population is sufficient to uniquely identify the mobility distribution. We reconstruct this distribution using a machine-learning-based framework, providing both theoretical and algorithmic support to effectively constrain the mobility-based model with real-world data.
@@ 1 Introduction
Understanding and predicting epidemic dynamics are crucial for public health planning and response [14, 9, 16]. Accurate models facilitate the timely implementation of control measures, resource allocation, and strategy development, which are essential for minimizing the impact of infectious diseases. For instance, during the COVID-19 pandemic, models by Ferguson et al. [7] guided decisions on lockdowns and social distancing, significantly influencing policy and saving lives.
Compartmental models have been a cornerstone in epidemiology since the early 20th century. These models, such as the classic SIR (Susceptible-Infected-Recovered) model introduced by Kermack and McKendrick [11], divide the population into compartments based on disease status. The transitions between compartments are governed by differential equations that describe the dynamics of disease transmission and recovery. These models have been instrumental in understanding various infectious diseases, including influenza, measles, and more recently, COVID-19. Compartmental models provide a simplified yet powerful framework to study the spread of diseases within a population.
Traditional compartmental models, despite their utility, face significant limitations [10, 1]. One major constraint lies in their assumption of a homogeneous mixing population, where every individual has an equal probability of coming into contact with any other individual, disregarding the complex realities of social interactions and geographic distributions [2, 15]. Furthermore, these models typically omit spatial dynamics, which is critical for understanding disease spread across different regions and the localized impact of interventions [6]. Additionally, traditional models often fail to account for dynamic changes in human behavior during epidemics, such as increased handwashing, social distancing, and vaccination uptake, which can profoundly influence epidemic outcomes [5, 8]. This oversimplification neglects the diverse patterns of human contact within communities.
The current ODE-based epidemiological models frequently overestimate pandemic size by assuming homogeneous mixing and neglecting real-world contact patterns [12, 13, 3]. Moreover, the classic models often fail to incorporate behavioral changes such as increased hygiene practices or timely interventions, which can inflate estimates of susceptible populations and contribute to the overall overestimation. To address these issues, we propose mobility-based compartmental models where the population exhibits heterogeneous mobility distributions. Individuals with higher mobility are at increased risk of infection and transmission during the early stages of the pandemic, potentially leading to a higher basic reproduction number ($\mathcal{R}_{0}$) initially. However, as the pandemic progresses, the initially infected high-mobility group recovers and subsequently provides indirect protection to the low-mobility group, thereby reducing $\mathcal{R}_{0}$ and ultimately limiting the pandemic size. This mobility-based infection mechanism offers a straightforward yet effective approach to more accurately estimate pandemic spread, addressing the overestimation observed in traditional compartmental models used in epidemiology.
@@
Our paper proceeds as follows. In section 2, we propose a mobility-based heterogeneous compartmental model, highlighting its distinctions from homogeneous models in terms of $\mathcal{R}_{0}$ and the final pandemic size. In section 3, we present an inverse problem aimed at inferring the mobility distribution from the time series of the infected population. We provide sufficient conditions to infer the mobility distribution uniquely and propose a machine-learning approach for inferring the mobility distribution. We conclude in section 4.
2 Model description and properties
@@ 2.1 A mobility-based SIRS model
The baseline SIRS model delineates infectious disease dynamics within a population, categorizing individuals into susceptible (S), infected (I), and recovered (R) compartments. Consider a system with $N$ agents, in which agent $i$ holds a state $X_{i}$ that takes a value S, I, or R. We define $S(t)$, $I(t)$, and $R(t)$ as the ratios of agents in the three states as
$S(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{S}\}},\quad I(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{I}\}},\quad R(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{R}\}}.$ (1)
The baseline SIRS model posits that individuals transition cyclically between S, I, and R states through infection, recovery, and subsequent loss of immunity, all modeled as independent Poisson-point processes with constant rates. The well-known SIRS model describes the population dynamics of the three compartments as
$\dot{S}(t)$ $=-{\beta}S(t)I(t)+\varepsilon R(t)$ (2)
$\dot{I}(t)$ $={\beta}S(t)I(t)-\gamma I(t)$
@@
However, this baseline model overlooks the heterogeneity among individuals within a population, which is critical for understanding disease transmission. Among various factors, mobility plays a pivotal role in epidemics by accelerating infection movement across regions through travel, commuting, and migration. We propose a mobility-based SIRS model where each agent $i$ has a time-independent mobility $m_{i}\in(0,1)$ that determines the infection and transmission rates of agent $i$. All S-I pairs (e.g., a pair of agents $i$ and $j$ with $X_{i}=\text{S}$ and $X_{j}=\text{I}$) are subject to internal infections, with the probability of agent $j$ infecting agent $i$ given by $\beta m_{i}m_{j}$. Infected agents recover at a constant rate $\gamma$, and recovered agents lose immunity at a constant rate $\varepsilon$, consistent with the baseline model (2). Figure 1 illustrates the transitions between the three compartments and their respective rates.
@@
We describe the dynamics using three population density functions of mobility—$S(m,t)$, $I(m,t)$, $R(m,t)$—associated with the susceptible, infected, and recovered compartments, respectively. Here, $S(m,t)\mathrm{d}m,$ represents the ratio of susceptible agents with mobility in the interval $[m,m+\mathrm{d}m,)$. Due to population conservation, we have
$S(m,t)+I(m,t)+R(m,t)=f(m),\quad\text{for all}\leavevmode\nobreak\ t\geq 0,$ (3)
where $f(m)$ denotes the time-independent mobility distribution function. The corresponding mobility-based SIRS model is
$\frac{\partial}{\partial t}S(m,t)$ $=-{\beta}\int_{0}^{1}mm^{\prime}S(m,t)I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},+\varepsilon R(m,t)$ (4a)
$\frac{\partial}{\partial t}I(m,t)$ $={\beta}\int_{0}^{1}mm^{\prime}S(m,t)I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},-\gamma I(m,t)$ (4b)
@@
In the mobility-based SIRS model, a susceptible agent with mobility $m$ can be infected by any infected agent with mobility $m^{\prime}$, leading to an integral representation of the internal infection process. In our model, each agent’s mobility is time-independent, resulting in local terms for the recovery and loss of immunity processes.
@@ 2.2 Dynamics of the mobility-based SIRS model
In Figure 2, we illustrate the time evolution of the three population densities in the mobility-based SIRS model (4) during the first infection wave. Initially, the population is evenly distributed across the mobility domain $[0,1]$. As the simulation begins, susceptible agents with high mobility are infected first, leading to a peak in the infected population density near $m=1$ at $t=250$. Over time, the infected population recovers, resulting in an increased density in $R(m,t)$. This infection-recovery process occurs earlier in the high-mobility region compared to the low-mobility region.
@@
Figure 2: Evolution dynamics of the mobility-based SIRS model (3), with parameter $\beta=0.1,\varepsilon=0.01$, $\gamma=0.3$ and initial $S(m,0)=0.99,I(m,0)=0.01,R(m,0)=0$, for (a) the susceptible population $S(m,t)$, (b) the infected population $I(m,t)$, and (c) the recovered population $R(m,t)$.
@@
and illustrate the time evolution of the ratios of three compartments in Figure 3. We fix the second-order moment of the initial mobility distribution (i.e., $\langle m^{2}f\rangle$) in all simulations, which, as we will show in section 2.3, is proportional to the basic reproduction number $\mathcal{R}_{0}$. In addition, we use the initial conditions
@@
which we refer to as the proportional initial condition. We will use this initial condition by default without further explanation. We choose three initial mobility distributions that all have the same second-order moment $\langle m^{2}f\rangle$, where $f_{1}$ and $f_{2}$ are restricted normal distributions on $[0,1]$ with
@@
and $f_{3}$ is a Direct delta function $f(m)=\delta(m-0.435)$.
@@ 2.3 Basic reproduction number
The basic reproduction number $\mathcal{R}_{0}$ is a key epidemiological metric that represents the average number of secondary infections produced by a single infected individual in a fully susceptible population. In this section, we derive $\mathcal{R}_{0}$ of the mobility-based SIRS model (4).
We consider the next-generation matrix method [4] and discretize $I(m,t)$ and $S(m,t)$ on the grids $\{m_{i}\}_{i=1}^{M}$. We denote the discretized quantities by two vectors $\bm{x},\bm{y}\in\mathbb{R}^{M}$, in which
$x_{i}(t)=I(m_{i},t),\quad y_{i}(t)=S(m_{i},t).$ (8)
We consider a numerical quadrature to replace the integral in (4) with
$\int_{0}^{1}m^{\prime}I(m^{\prime},t)\leavevmode\nobreak\ \mathrm{d}m^{\prime},\approx\sum_{i=1}^{M}w_{i}m_{i}x_{i},$ (9)
where $w_{i}$ is the weight of the numerical quadrature at the grid point $m_{i}$. We write (4) in a discrete format and obtain
$\dot{\bm{x}}=\bm{F}(\bm{x},\bm{y})-\bm{V}(\bm{x},\bm{y}),$ (10)
where $\bm{F}$ is the rate of new infections from the susceptible population and $\bm{V}$ is the rate of loss of infections due to recovery. They are
$F_{i}=\beta m_{i}y_{i}\sum_{j}w_{j}m_{j}x_{j},\quad V_{i}=\gamma x_{i}.$ (11)
Consider the disease-free equilibrium, in which the whole population only contains susceptible individuals and its population density satisfies $S(m,t)\equiv f(m)$. Therefore, the disease-free equilibrium vectors $\bm{x}^{0},\bm{y}^{0}\in\mathbb{R}^{M}$ satisfy $x^{0}_{i}=f(m_{i})$ and $y^{0}_{i}=0$. We compute the Jacobi matrices of $\bm{F}$ and $\bm{V}$ with respect to $\bm{x}$ at the disease-free equilibrium vectors $\bm{x}^{0},\bm{y}^{0}$ and obtain
$\frac{\partial\bm{F}}{\partial\bm{x}}(\bm{x}^{0},\bm{y}^{0})_{ij}=\beta m_{i}y^{0}_{i}m_{j}w_{j},\quad\frac{\partial\bm{V}}{\partial\bm{x}}(\bm{x}^{0},\bm{y}^{0})_{ij}=\gamma\delta_{ij},$ (12)
where $\delta_{ij}$ is the Kronecker delta function. We further compute the next-generation matrix
$\frac{\partial\bm{F}}{\partial\bm{x}}\left(\frac{\partial\bm{V}}{\partial\bm{x}}\right)^{-1}=\frac{\beta}{\gamma}\begin{pmatrix}m_{1}y^{0}_{1}\\ \vdots\\ m_{M}y^{0}_{M}\end{pmatrix}\begin{pmatrix}m_{1}w_{1}&\cdots&m_{M}w_{M}\end{pmatrix}_{j}.$ (13)
We compute the largest eigenvalue of the next-generation matrix and obtain the basic reproduction number $\mathcal{R}_{0}=\beta\sum_{i}w_{i}m_{i}^{2}y^{0}_{i}/\gamma$. Notice that $w_{i}$ are the weights of grid points in the numerical quadrature, so as the number of grids $M$ goes to infinite, $\mathcal{R}_{0}$ converges to the integral
@@
$\mathcal{R}_{0}=\frac{\beta}{\gamma}\int_{0}^{1}m^{2}f(m)\leavevmode\nobreak\ \mathrm{d}m,.$ (14)
@@
The above derivations can be made rigorous by using the Fréchet derivative with an appropriate definition of the function space. When the entire population has a constant mobility $m_{0}$, equivalent to the baseline SIRS model (2), the basic reproduction number in (14) is
@@
which is consistent with the results of the baseline SIRS model (2).
Let $\text{var}(f)$ denote the variance of a mobility distribution $f$. Assuming that all distributions that have the same $\mathcal{R}_{0}$, from (14) we have
@@
where $f_{\delta}$ is the Dirac delta distribution that has the same second-order moment as other distributions. Therefore, if we fix $\mathcal{R}_{0}$, the Dirac delta distribution $f_{\delta}$ maximizes the mean mobility $\langle mf\rangle$. In practice, we often estimate $\mathcal{R}_{0}$ by considering the early growth rate of cases and performing contact tracing studies, which is independent of model selection. As a result, using the value of $\mathcal{R}_{0}$ to estimate the pandemic size with the baseline SIRS model tends to result in overestimation.

### v2

@@ Abstract
Classical compartmental models in epidemiology often assume a homogeneous population for simplicity, which neglects the inherent heterogeneity among individuals. This assumption frequently leads to inaccurate predictions when applied to real-world data. For example, evidence has shown that classical models overestimate the final pandemic size in the H1N1-2009 and COVID-19 outbreaks. To address this issue, we introduce individual mobility as a key factor in disease transmission and control. We characterize disease dynamics using mobility distribution functions for each compartment and propose a mobility-based compartmental model that incorporates population heterogeneity. Our results demonstrate that, for the same basic reproduction number, our mobility-based model predicts a smaller final pandemic size compared to the classical models, effectively addressing the common overestimation problem. Additionally, we infer mobility distributions from the time series of the infected population. We provide sufficient conditions for uniquely identifying the mobility distribution from a dataset and propose a machine-learning-based approach to learn mobility from both synthesized and real-world data.
@@ 1 Introduction
Understanding and predicting epidemic dynamics are essential for effective public health planning and response [23, 14, 26]. Reliable and effective models provide crucial insights that enable timely implementation of control measures, efficient resource allocation, and strategic planning. For example, during the COVID-19 pandemic, models by Ferguson et al. [11] and IHME [5] guided decisions on school closures, social distancing, mask mandates, and reopening. Epidemiological models significantly influenced policy and played a key role in saving lives.
Among all these models, compartmental models have served as a cornerstone in epidemiology since the early 20th century. These compartmental models, including the well-known SIR (Susceptible-Infected-Recovered) model introduced by Kermack and McKendrick [18] in 1927, divide the individuals into compartments based on their disease status and use ordinary differential equations to describe the dynamic evolution of each compartment. Despite their utility, classical compartmental models have significant limitations [17, 1]. A major issue is their assumption of a homogeneous mixing population, where each individual has an equal chance of interacting with any other. This overlooks spatial dynamics [10, 28], real-world contact patterns [20, 21, 3], and dynamic changes in human behavior during epidemics [9, 12], such as increased handwashing, social distancing, and vaccination, which can significantly affect epidemic outcomes. The COVID-19 pandemic further highlights that classical models often ignore behavioral changes and interventions, leading to inflated estimates of susceptible populations.
To address these limitations, we propose mobility-based compartmental models incorporating heterogeneous mobility distributions. These models describe the time evolution of population densities for each compartment as a function of a mobility variable ranging from $[0,1]$, making them infinite-dimensional. While the literature has explored similar infinite-dimensional models for their basic mathematical properties, to the best of our knowledge, this paper addresses critical topics in epidemiology such as the estimation of basic reproduction number and final pandemic size and the identifiability of the mobility distribution, which have not been previously studied.
A notable innovation in our paper is the introduction of individual mobility, a key factor in infectious disease spread. During an outbreak, individuals with higher mobility face greater risks of infection and transmission early on. As the pandemic progresses, this high-risk group recovers and indirectly protects the low-mobility group, thereby slowing disease spread and limiting the overall pandemic size. Classical ODE-based compartmental models cannot capture this mechanism. We prove that for the same basic reproduction number $\mathcal{R}_{0}$, a mobility-based SIR model predicts a smaller final pandemic size compared to a classical SIR model, addressing the common overestimation issue in the literature.
Another key contribution of this paper is the inference of mobility distribution from real-life data. For accurate forecasts and to match observed data, it is crucial to use a mobility distribution from real-world measurements. However, obtaining reliable mobility data is challenging due to large population sizes, privacy constraints, and inherent variability in individual movement patterns. We propose a method to infer the mobility distribution from accessible pandemic data, specifically the ratio of the infected population $I(t)$. We prove a one-to-one correspondence between the mobility distribution and the time series of $I(t)$, implying that the inference problem admits a unique solution. We also construct a training set of mobility distributions using Gaussian mixtures and employ an artificial-neural-network approach to identify the mobility distribution from real-life data.
@@
Our paper proceeds as follows. In section 2, we introduce a mobility-based compartmental model and emphasize its distinctions from homogeneous models, particularly in terms of the basic reproduction number and the final pandemic size. In section 3, we formulate the inverse problem to infer the mobility distribution from the time series of the infected population. We prove this inverse problem has a unique solution and propose a machine-learning approach to infer the mobility distribution. We conclude in section 4.
2 Model description and epidemiological parameters
@@ 2.1 A mobility-based SIRS model
The baseline SIRS model categorizes agents into three compartments: susceptible (S), infected (I), and recovered (R). Consider a system with $N$ agents, in which agent $i$ holds a disease state $X_{i}$ that takes a value of S, I, or R. We define $S(t)$, $I(t)$, and $R(t)$ as the ratios of agents in the three states:
$S(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{S}\}},\quad I(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{I}\}},\quad R(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{1}_{\{X_{i}(t)=\text{R}\}},$ (1)
where $\mathbf{1}$ is the indicator function. The baseline SIRS model assumes individuals cycling between the S, I, and R states through infection, recovery, and loss of immunity. The baseline SIRS model without birth and death reads as
$\dot{S}(t)$ $=-{\beta}S(t)I(t)+\varepsilon R(t),$ (2)
$\dot{I}(t)$ $={\beta}S(t)I(t)-\gamma I(t),$
@@
Mobility is a key factor in epidemics, influencing the spread of infection across regions through travel, commuting, and migration. We let each agent have a time-independent mobility $m_{i}\in(0,1)$, which determines their infection and transmission rates. For any S-I pair (i.e., a susceptible agent $i$ and an infected agent $j$), the probability of infection is given by $\beta m_{i}m_{j}$. Infected agents recover at a constant rate $\gamma$, and recovered agents lose immunity at a constant rate $\varepsilon$. Figure 1 illustrates the transitions between the three compartments and their respective rates.
@@
We describe the dynamics using three population density functions of mobility—$S(m,t)$, $I(m,t)$, $R(m,t)$—associated with the susceptible, infected, and recovered compartments, respectively. Here, $S(m,t)\mathrm{d}m$ represents the ratio of susceptible agents with mobility in the interval $[m,m+\mathrm{d}m)$. Let $f$ be the mobility distribution function of the whole population and we assume $f$ is time-independent. Due to population conservation, we have
$S(m,t)+I(m,t)+R(m,t)=f(m),\quad\text{for all}\leavevmode\nobreak\ t\geq 0.$ (3)
While individuals with different mobility can interact in numerous ways, we assume that interactions between susceptible and infected individuals still follow the mass-action principle, as in the classical SIR model. We assume that the probability of infection between an S-I pair is proportional to the product of their mobilities. Therefore, the corresponding mobility-based SIRS model is
$\frac{\partial}{\partial t}S(m,t)$ $=-{\beta}\int_{0}^{1}m\bar{m}S(m,t)I(\bar{m},t)\leavevmode\nobreak\ \mathrm{d}\bar{m}+\varepsilon R(m,t),$ (4a)
$\frac{\partial}{\partial t}I(m,t)$ $={\beta}\int_{0}^{1}m\bar{m}S(m,t)I(\bar{m},t)\leavevmode\nobreak\ \mathrm{d}\bar{m}-\gamma I(m,t),$ (4b)
@@
In the mobility-based SIRS model, a susceptible agent with mobility $m$ can be infected by any infected agent with mobility $\bar{m}$, leading to an integral representation of the internal infection process. In our model, each agent’s mobility is time-independent, resulting in local terms for the recovery and loss of immunity processes.
@@ 2.2 Dynamics of the mobility-based SIRS model
In Figure 2, we illustrate the time evolution of the mobility-based SIRS model (4) during the first infection wave. Initially, the population is evenly distributed across the mobility domain $[0,1]$. As the simulation begins, susceptible agents with high mobility are infected first, leading to a peak in the infected population density near $m=1$ at $t=250$. Over time, the infected population recovers, resulting in an increased density in $R(m,t)$. This infection-recovery process occurs sooner in the high-mobility region compared to the low-mobility region.
@@
Figure 2: Evolution dynamics of the mobility-based SIRS model (3), with parameter $\beta=0.1,\varepsilon=0.01$, $\gamma=0.3$ and initial $S(m,0)=0.99,I(m,0)=0.01,R(m,0)=0$ for (a) the susceptible population $S(m,t)$, (b) the infected population $I(m,t)$, and (c) the recovered population $R(m,t)$.
@@
and illustrate the time evolution of the ratios of three compartments in Figure 3. We use the initial conditions
@@
which we refer to as the proportional initial condition. To demonstrate the effect of heterogeneous mobility, we choose three initial mobility distributions $f_{1}$, $f_{2}$, and $f_{3}$, where $f_{1}$ and $f_{2}$ are restricted normal distributions on $[0,1]$ with
@@
and $f_{3}$ is a Direct delta function $f_{2}(m)=\delta(m-0.435)$. All three mobility distributions have the same second-order moment $\langle m^{2}f\rangle$, which, as we will show in section 2.3, is proportional to the basic reproduction number. The mobility distribution $f_{1}$ is a mixture of two restricted Gaussian distributions and has two bumps centered at $m=0.1$ and $m=0.95$, indicating a polarized distribution of mobility among the population. The mobility distribution $f_{3}$ corresponds to the baseline SIRS model (2) in which everyone has the same mobility.
@@
We observe from Figure 3 that at the end of the pandemic, over $70\%$ of the population are infected for the homogeneous population case $f_{3}$. In contrast, $f_{1}$ has polarized mobility distribution with a relatively small group of high-mobile individuals. When high-mobile individuals recover from infection, they can provide protection to uninfected low-mobile groups. Figure 3 shows that the final pandemic size regarding $f_{1}$ is only about $1/3$ of that of the baseline model. We derive and discuss the final pandemic size in detail in section 2.5.
@@ 2.3 Basic reproduction number
The basic reproduction number $\mathcal{R}_{0}$ is a key epidemiological metric that represents the average number of secondary infections produced by a single infected individual in a fully susceptible population. In this section, we derive the basic reproduction number $\mathcal{R}_{0}$ of the mobility-based SIRS model (4) by using the next-generation operator (i.e., threshold operator) framework [25], which is essentially the infinite-dimensional version of the next generation matrix method [8].
Recall the definition of $\mathcal{R}_{0}$ in [25]. For a generic infinite-dimensional SIRS model, suppose the infected population is defined on a compact set $\Omega$ and satisfies
$\frac{\partial I(x,t)}{\partial t}=\int_{\Omega}h(x,y,S(x,t),I(y,t))\leavevmode\nobreak\ \mathrm{d}y-v(x)I(x,t),$ (8)
where $h(x,y,S,I)$ is an incidence function and $v(x)$ is the recovery rate function. We define the threshold operator $\mathcal{L}$ as
@@
$\mathcal{L}u(x)=\int_{\Omega}\frac{u(y)}{v(x)}\left.\frac{\partial h(x,y,S,I)}{\partial I}\right|_{S=N(x),I=0}\leavevmode\nobreak\ \mathrm{d}y,$ (9)
where $N(x)$ is the total population density, then the basic reproduction number $\mathcal{R}_{0}$ is the spectral radius of $\mathcal{L}$, which means
$\mathcal{R}_{0}=\lim_{n\rightarrow\infty}\|(\mathcal{L})^{n}\|^{1/n}.$ (10)
We refer [25] for detailed assumptions on $h$ and $v$.
For the mobility-based SIRS model defined in (4), the total population density is $f(m)$ and the threshold operator $\mathcal{L}^{f}$ is
$\mathcal{L}^{f}u(m)=\frac{\beta}{\gamma}mf(m)\int_{0}^{1}\bar{m}u(\bar{m})\mathrm{d}\bar{m}.$ (11)
Therefore, the basic reproduction number $\mathcal{R}_{0}=\lim_{n\rightarrow\infty}\|(\mathcal{L}^{f})^{n}\|^{1/n}$, and we compute $\mathcal{R}_{0}$ in Theorem 1.
Theorem 1.
The basic reproduction number $\mathcal{R}_{0}$ of the mobility-based SIRS model (4) is
$\mathcal{R}_{0}=\frac{\beta}{\gamma}\langle m^{2}f\rangle=\frac{\beta}{\gamma}\int_{0}^{1}m^{2}f(m)\leavevmode\nobreak\ \mathrm{d}m.$ (12)
Proof.
Recall the definition of the threshold operator $\mathcal{L}^{f}$ in (11). For any function $u(m)$, the image $\mathcal{L}^{f}u(m)$ lies in a 1-dimensional function space spanned by $mf(m)$. Consequently, $\mathcal{L}^{f}$ is a finite-rank operator, and thus, it must be compact. Therefore, the spectral radius of $\mathcal{L}^{f}$ is its maximal eigenvalue.
Because $\mathcal{L}^{f}$ is a rank-one operator, $\mathcal{L}^{f}$ only has 1 nonzero eigenvalue and this nonzero eigenvalue corresponds to the eigenfunction $mf(m)$. We compute this eigenvalue by applying $\mathcal{L}^{f}$ to $mf(m)$ and obtaining
$\mathcal{L}^{f}(mf(m))=\frac{\beta}{\gamma}mf(m)\int_{0}^{1}\bar{m}^{2}f(\bar{m})\mathrm{d}\bar{m}=\frac{\beta}{\gamma}\langle m^{2}f\rangle mf(m).$ (13)
Therefore, the leading eigenvalue of $\mathcal{L}^{f}$ is $\frac{\beta}{\gamma}\langle m^{2}f\rangle$ and it is also the basic reproduction number $\mathcal{R}_{0}$.
The basic reproduction number, $\mathcal{R}_{0}$, is a key dimensionless epidemiological parameter that represents the average number of secondary cases generated by one primary case. When $\mathcal{R}_{0}>1$, the disease-free equilibrium becomes unstable, allowing the disease to persist in the population. Conversely, if $\mathcal{R}_{0}<1$, the disease-free equilibrium is asymptotically stable, leading to the eventual extinction of the disease. The stability of both disease-free and endemic equilibria has been extensively studied over the past decades. For the original next-generation matrix method applied to finite-dimensional compartmental models, see [8, 7], and for various infinite-dimensional generalizations, refer to [24, 25, 22]. By applying Theorems 3.16 and 3.17 in [24], one can readily show that the disease-free equilibrium $(f(m),0,0)$ of our model (4) is unstable when $\mathcal{R}_{0}>1$ and asymptotically stable when $\mathcal{R}_{0}<1$.
@@
When the entire population has a constant mobility $m_{0}$, equivalent to the baseline SIRS model (2), the basic reproduction number in (12) becomes
@@
which is the same as the basic reproduction number of the baseline SIRS model (2) with the infection rate $\beta_{\text{eff}}$.
Let $\text{var}(f)$ denote the variance of a mobility distribution $f$. In practice, we often estimate $\mathcal{R}_{0}$ by considering the early growth rate of cases and performing contact tracing studies, which is independent of models. Assuming that all distributions that have the same $\mathcal{R}_{0}$, from (12) we have
@@
where $f_{\delta}$ is the Dirac delta distribution that has the same second-order moment as $f$. Therefore, the mean mobility $\langle mf_{\delta}\rangle$ of the Dirac delta distribution $f_{\delta}$ is larger than that of any other distribution $f$. Once $\mathcal{R}_{0}$ is determined first (using a model-free method), the homogeneous model (e.g., the baseline SIRS model) overestimates the mean mobility, which consequently leads to an overestimation of the final pandemic size.